# Intermittent localization and fast spatial learning by non-Markov random walks with decaying memory

### v1

@@ Abstract
Random walks on lattices with preferential relocation to previously visited sites provide a simple modeling of the displacements of animals and humans. When the lattice contains a single impurity or resource site where the walker spends more time on average at each visit than on the other sites, the long range memory can suppress diffusion and induce by reinforcement a steady state localized around the resource. This phenomenon can be identified with a spatial learning process by the walker. Here we study theoretically and numerically how the decay of memory impacts learning in these models. If memory decays as $1/\tau$ or slower, where $\tau$ is the time backward into the past, the localized solutions are the same as with perfect, non-decaying memory and they are linearly stable. If forgetting is faster than $1/\tau$, for instance exponential, an unusual regime of intermittent localization is observed, where well localized periods of exponentially distributed duration are interspersed with intervals of diffusive motion. At the transition between the two regimes, for a kernel in $1/\tau$, the approach to the stable localized state is the fastest, opposite to the expected critical slowing down effect. Hence forgetting can allow the walker to save memory without compromising learning and to achieve a faster learning. These findings agree with biological evidence on the benefits of forgetting.
@@ I Introduction
Whereas the lattice random walk represents a paradigm for studying diffusion, some random walk models exhibit localization properties that are in or out-of-equilibrium [1, 2, 3, 4]. A classical example is the so-called vertex reinforced random walk, consisting of a walker that hops to nearest-neighbor sites, such that the probability to transit to a site depends linearly on the number of previous visits received by that site in the past [5, 6, 7]. As the walker needs to keep track of all the visits to all the sites, the process has unbounded memory and is highly non-Markov unlike the ordinary random walk. A consequence of local attraction toward more frequently visited sites can be the suppression of diffusion and the emergence of localization. The reinforced random walk in one dimension ($1d$), for instance, exhibits a highly non-trivial property: it gets stuck asymptotically with probability 1 on five sites of the lattice, that are visited infinitely often [8].
@@
Other models based on the principle of preferential returns to previously visited locations have been developed over the two past decades and compared quantitatively to real movement data of single animals [9, 10, 11, 12] and humans [13]. Empirical observations actually indicate that individuals across many species use their memory and tend to return preferentially to familiar places in their environment [14], i.e., that the probability to revisit a location is roughly proportional to the total amount of time spent at this location previously [13, 10]. It is also well established that most animals have home ranges, i.e., are localized on a rather limited area instead of diffusing randomly and unbounded. The mechanisms of home range formation are not completely understood [15], but memory is assumed to play a central role [16]. However, most ecological movement models based on preferential revisits, although capturing realistic features of field studies, do not commonly exhibit localization but rather very slow diffusion.
@@
An analytically solvable example is the preferential relocation model or ’monkey walk’ [10, 17, 18, 19], that mixes random exploration with self-attraction. At each discrete time-step, a walker on an infinite $1d$ lattice takes a usual random move to a nearest-neighbor with probability 1q1-q or resets to a previous site (say $j$) with the complementary probability $q$. In the latter eventuality, the site $j$ is chosen among all the visited sites with a probability proportional to the total amount of time spent there since the beginning of the process. One important difference with the classic reinforced random walk resides in the fact that the chosen site for revisit does not need to be a nearest neighbor of the walker’s current position. For any non-zero $q$, it is found that the occupation probability is always time-dependent and obeys a Gaussian scaling-law at large times, but with a variance growing extremely slowly, as $\ln(t)$. In other words, the process does not converge toward a stationary state (that would indicate localization), but evolves much more slowly than in normal diffusion, where the mean square displacement grows as $t$.
@@
A simple modification of the above model does exhibit localization, though. Assume that one particular site, placed at the origin, is more attractive than the other ones: each time the walker occupies the origin, it stays there for one more time-step with probability $\gamma$ and does the monkey walk with probability $1-\gamma$. On the other sites, the walker only performs monkey walk moves. This setup mimics an animal that feeds occasionally on a particular resource site, before continuing its motion in a scarce environment. In the limit $t\to\infty$, an analytical calculation shows that, in $1d$, a localized solution independent of time exists for any non-zero values of the probabilities $q$ and $\gamma$ [20]. The steady state occupation probability is centered around the resource site and is no longer Gaussian but with exponential tails, which define a finite localization length. In other cases, the walk is localized only if $q>q_{c}$ with $q_{c}>0$. Overall, the description of this phase transition is very similar phenomenologically to the self-consistent theory of Anderson localization, see [20, 21]. Extensions of this model to several resource sites exhibit similar features, with an enhancement of localization on the stronger impurities [22].
@@
There exists a fundamental difference between this type of localization and the ones observed in the classic reinforced random walk, or in the Markov random walk with resetting to the starting position only [23, 24, 25, 26, 27, 28, 29, 30]: the steady state is independent of the initial position and represents an attractor of the dynamics. This attractor is largely determined by the medium itself (the position and attractiveness $\gamma$ of the resource site) and results from the experience of the walker accumulated over time. At late times, the walker settles down around the impurity, in a way actually very similar to a memory-less random walker with resetting to the impurity only [20, 21], but without knowing before hand the existence of this convenient resetting position on the lattice. Hence the convergence toward localization can be identified as a learning process, in the sense understood in the behavioral sciences, i.e., a change in behavior resulting from experience [31].
@@
The purpose of the present study is to analyze the effects of a decaying memory on the existence of localized solutions and on their stability, in the learning monkey walk problem described above. Random walk models with preferential revisits usually assume that memory does not decay with time, or all visits are counted equally. Assuming that memory is not perfect, e.g., that more recent visits are remembered better than those of the remote past, is more realistic biologically. Exponential or power-law memory kernels have been identified by fitting animal ranging data to models with memory decay [32, 33, 11]. In preferential relocation processes, the main effects of memory decay is a possible modification, on homogeneous lattices, of the logarithmic diffusion law to a faster one in $t^{\mu}$ (with $0<\mu\leq 1$) [34, 17, 35], or changes in their relaxation dynamics under confinement [36].
@@
In the neurosciences, forgetting, besides being unavoidable, does not always represent a dysfunction of the brain but has some virtues: it is thought to allows to suppress interferences between different memories and to make room for retaining new information [37, 38, 39]. From the modeling side, mean-field approaches show that exponential forgetting can actually help foraging animals to cope with changing environments [40]. In numerical simulations of a swarm of interacting monkey walks in a static but heterogeneous environment, the decay of memory was seen to improve the aggregation of individuals around the best resource sites, a phenomenon identified as a successful collective learning [41]. Overall, very few analytical results are available on the effects of forgetting on spatial learning. To start with, is perfect memory a necessary condition for localization in preferential random walk models?
@@
We now present the structure of the article and a summary of the main results. In Section II we introduce the memory random walk model with one impurity on the infinite lattice, where forgetting is incorporated via a general kernel. In Section III, we recall known analytical results on the steady state solution of the problem with perfect memory. Section IV reports simulation results. When memory has an asymptotic decay slower than $1/\tau$ (where $\tau$ is the time backward into the past), i.e., of the form $1/\tau^{\beta}$ with $0\leq\beta\leq 1$, we find that the system tends toward the same localized state at late times as with perfect memory or $\beta=0$ (Section IV.1). This fact is consistent with the prediction obtained from a decoupling approximation. However, with kernels decaying faster than $1/\tau$ (Sections IV.2 and IV.3), an unusual dynamical behavior is observed: localization becomes intermittent in time, or spatial learning is disrupted, i.e., well-localized periods of finite durations are separated by intervals of diffusive motion. Somehow surprisingly, during these localized periods, the occupation probability is very close to the solution with perfect memory. In Section V, we perform a linear stability analysis of the fully localized solution using the decoupling approximation, which turns out to be consistent only for memory kernels decaying more slowly than $1/\tau$. In these cases, we find that the localized solutions are stable, in agreement with the numerics of Section IV.1. We then study the relaxation at late times toward these stable states (Section VI), showing that it is non-exponential and of the form $t^{\nu}/\ln t$, with $\nu\in(-1,0)$ a non-trivial exponent solution of Eq. (53). In the boundary case $\beta=1$ corresponding to the $1/\tau$ memory kernel at the edge of intermittency, the localized state is still stable and the relaxation toward the latter is actually the fastest of all the values of $\beta\in[0,1]$: the amplitudes of the perturbations tend to zero no more slowly than $(\ln t)^{\alpha}/t$, with $\alpha$ an exponent [Eqs. (58)-(59)]. This speed-up, confirmed by numerical simulations, is opposite to the usual critical slowing down for a state that is marginally stable. Hence the $1/\tau$ forgetting law is in some sense optimal, as $i)$ it achieves the same localization as perfect memory by being the most economical on memory, and ii)ii) it localizes the fastest. In Section VII we discuss different possible scenarios for intermittent localization (or imperfect learning) and conclude in Section VIII.
@@ II The model
We revisit the discrete time random walk model with preferential relocation to previously visited sites, in the presence of one impurity [20]. The original model considers a $1d$ lattice and a random walker with perfect memory that performs symmetric nearest-neighbor random steps with probability 1q1-q, and resets to a previously visited site with probability $q$. In the latter case, a site is chosen, among the visited sites, with a probability proportional to the total amount of time spent on that site since $t=0$. Additionally, there is an impurity site located at the origin $n=0$, representing a resource site with food. When the walker, with position $X_{t}$ at time $t$, is located at the origin (or $X_{t}=0$), it stays on that site at the next time-step with a probability $\gamma$ and moves according to the two movement rules above with the complementary probability $1-\gamma$. Hence, the parameter $0\leq\gamma\leq 1$ represents the attractiveness of the origin, which is occupied for a longer time at each visit compared to the other sites of the lattice. This site is likely to be more strongly reinforced by the resetting dynamics.
As mentioned above, when memory is used during the time step $t\to t+1$, a previous site is chosen to be revisited with a probability proportional to its accumulated amount of time. This linear rule is equivalent to choosing a time $t^{\prime}$ in the past, $t^{\prime}=0,1,...,t$, with uniform probability and to relocate to the position occupied at time $t^{\prime}$, i.e., to set
@@
A site which has been occupied often will have more possibilities (times $t^{\prime}$) to be chosen. The fact that $t^{\prime}$ is distributed uniformly also means that the walker remembers equally well all its previous positions or memory does not decay.
@@
Let us denote as $P_{n}(t)$ the occupation probability of the site $n$ by the walker at time $t$, $P_{n}(t)={\rm Prob}[X_{t}=n]$, given an unspecified initial condition. This quantity satisfies the following master equation,
@@
The first line of Eq. (II) represents the Markov random walk steps and trapping by the impurity, where $\delta_{i,j}$ is the Kronecker function. The second line accounts for the probability of revisiting the site $n$ via memory at time $t+1$ (because it was visited at a previous time $t^{\prime}$) by jumping from the origin which is occupied at time $t$. The third line counts the relocation jumps to $n$ from another site than the origin. The key point in this problem is that the resetting probability actually depends on the position: it is $q(1-\gamma)$ on the origin (since the walker does not move with probability $\gamma$) and $q$ on the other sites. Such heterogeneity is crucial for the existence of localized solutions, but it also makes the problem very difficult to solve, as Eq. (II) cannot be reduced to a closed form involving the marginal $P_{n}(t)$ only. Clearly, the master equation explicitly depends on higher-order joint probabilities such as ${\rm Prob}[X_{t^{\prime}}=n\ {\rm and}\ X_{t}=0]$, therefore a whole hierarchy should be written for the multiple-time distribution functions [21].
@@ III Previous results for the perfect memory case
We summarize the results obtained from the uniform memory kernel $\pi_{t,t^{\prime}}=1/(t+1)$ corresponding to no memory decay [20]. One first assume that $P_{n}(t)$ tends to a stationary distribution $P_{n}^{(st)}>0$ at large times for all $n$, where $P_{n}^{(st)}$ is peaked around the origin. This solution, if it exists, corresponds to a localized phase, by opposition to a late time-dependent scaling solution that would tend to 0 uniformly, of the form $P_{n}(t)\sim\phi(t)^{-1}G[n/\phi(t)]$ with $\phi(t)\to\infty$. The latter form with $\phi(t)\sim\ln t$ was found for the case $\gamma=0$ (absence of impurity), a problem which is more tractable since Eq. (II) closes in that case [10, 35, 17]. Some progress can be made in the resolution of Eq. (II) with $\gamma\neq 0$ by noticing that the times $t^{\prime}$ and $t$ are in general very far apart asymptotically, justifying the use of a de-correlation approximation,

### v3

@@ Abstract
Random walks on lattices with preferential relocation to previously visited sites provide a simple framework for modeling the displacements of animals and humans. When the lattice contains a few impurities or resource sites where the walker spends more time on average at each visit than on the other sites, the long range memory can suppress diffusion and induce by reinforcement a steady state localized around a resource. This phenomenon can be identified with a spatial learning process. Here we study theoretically and numerically how the decay of memory impacts learning in a model with one impurity. If memory decays as $1/\tau$ or slower, where $\tau$ is the time backward into the past, the localized solution is the same as with perfect, non-decaying memory and it is linearly stable. If forgetting is faster than $1/\tau$, for instance exponential, an unusual regime of intermittent localization is observed, where well localized periods of exponentially distributed duration are disrupted by possibly long intervals of diffusive motion. At the transition between the two regimes, for a kernel in $1/\tau$, the approach to the stable localized state is the fastest, opposite to the expected critical slowing down effect. Hence, forgetting can allow the walker to save a lot of memory without compromising learning and to achieve a faster learning process. These findings agree with biological evidence on the benefits of forgetting.
@@ I Introduction
Whereas the lattice random walk represents a paradigm for studying diffusion, some random walk models exhibit localization properties that are in or out-of-equilibrium [1, 2, 3, 4]. A classical example is the so-called vertex reinforced random walk, consisting of a walker that hops to nearest-neighbor (n.n.) sites, such that the probability of transiting to a site depends linearly on the number of previous visits received by that site in the past [5, 6, 7]. As the walker needs to keep track of all the visits to all the sites, the process has unbounded memory and is highly non-Markov unlike the ordinary random walk. A consequence of local attraction toward sites that are visited more frequently can be the suppression of diffusion and the emergence of localization. The reinforced random walk in one dimension ($1d$), for instance, exhibits a highly non-trivial property: it gets stuck asymptotically with probability 1 on five sites of the lattice, that are visited infinitely often [8].
@@
Other models based on the principle of preferential returns to previously visited locations have been developed over the past two decades and quantitatively compared to real movement data of single animals [9, 10, 11, 12] or humans [13]. Empirical observations actually indicate that individuals across many species use their memory and tend to return preferentially to familiar places in their environment [14], i.e., that the probability of revisiting a location is roughly proportional to the total amount of time spent at this location previously [13, 10]. It is also well established that most animals have home ranges, i.e., are localized on a rather limited area instead of diffusing randomly and unbounded. The mechanisms of home range formation are still not fully understood [15], but memory is assumed to play a central role [16]. However, most ecological movement models based on preferential revisits, although capturing realistic features of field studies, do not commonly exhibit localization but rather very slow diffusion.
@@
An analytically solvable example is the preferential relocation model or the ’monkey walk’ [10, 17, 18, 19], which mixes random exploration with self-attraction. At each discrete time-step, a walker on an infinite $1d$ lattice takes a usual random n.n. move with probability 1q1-q or resets to a previous site (say $j$) with the complementary probability $q$. In the latter eventuality, the site $j$ is chosen among all visited sites with a probability proportional to the total amount of time spent there since the beginning of the process. One important difference from the classic reinforced random walk lies in the fact that the chosen site for a revisit does not need to be a nearest neighbor of the walker’s current position. For any non-zero $q$, it is found that the occupation probability is always time-dependent and obeys a Gaussian scaling-law at large times, but with a variance growing extremely slowly, as $\ln(t)$. In other words, the process does not converge towards a stationary state (which would indicate localization), but evolves much more slowly than in normal diffusion, where the mean square displacement increases as $t$.
@@
A simple modification of the above model does exhibit localization, though. Assume that one particular site, placed at the origin, is more attractive than the other ones: each time the walker occupies the origin, it stays there for one more time-step with probability $\gamma$ and does the monkey walk with probability $1-\gamma$. On the other sites, the walker only performs monkey walk moves. This setup mimics an animal that occasionally feeds on a particular resource site before continuing its motion in a scarce environment. In the limit $t\to\infty$, analytical arguments show that, in $1d$, a localized solution independent of time exists for any non-zero values of the probabilities $q$ and $\gamma$ [20]. The steady state occupation probability is centered around the resource site and is no longer Gaussian but with exponential tails, which define a finite localization length. If the underlying Markov process between resetting events is not recurrent, e.g., the walker performs a $1d$ Lévy flight with an index smaller than unity, or a n.n. random walk in $3d$, the process localizes only if $q>q_{c}$ with $q_{c}>0$. Overall, the description of this phase transition is very similar phenomenologically to the self-consistent theory of Anderson localization [20, 21]. Extensions of this model to several resource sites exhibit similar features, including the enhancement of localization on the stronger impurities [22].
@@
There exists a fundamental difference between this type of localization and the ones observed in the classic reinforced random walk, or in the Markov random walk with resetting to the starting position only [23, 24, 25, 26, 27, 28, 29, 30]: the steady state is independent of the initial position and represents an attractor of the dynamics. This attractor is largely determined by the medium itself (the position and attractiveness $\gamma$ of the resource site) and results from the experience of the walker accumulated over time. At late times, the walker settles down around the impurity, in a way actually very similar to a memory-less random walker with resetting to the impurity only [20, 21], but without knowing before hand the existence of this convenient resetting position on the lattice. Hence, convergence toward localization can be identified as a learning process, in the sense understood in the behavioral sciences, i.e., a change in behavior resulting from experience [31].
@@
The purpose of the present study is to analyze the effects of a decaying memory on the existence of localized solutions and on their stability, in the learning monkey walk problem described above. Random walk models with preferential revisits usually assume that memory does not decay with time, or all visits are counted equally. Assuming that memory is not perfect, e.g., that more recent visits are remembered better than those of the remote past, is more realistic biologically. Exponential or power-law memory kernels have been identified, for instance, by fitting animal ranging data to models with memory decay [32, 33, 11]. In preferential relocation processes, the main effect of memory decay is a possible modification, on homogeneous lattices, of the logarithmic diffusion law to a faster one, in $t^{\mu}$ with $0<\mu\leq 1$ [34, 17, 35], or changes in their relaxation dynamics under confinement [36].
@@
In the neurosciences, forgetting, apart from being unavoidable, does not always represent a dysfunction of the brain but has some virtues: it contributes to suppress interferences between different memories and makes room for retaining new information [37, 38, 39]. From the modeling side, mean-field approaches show that exponential forgetting can actually help foraging animals to cope with changing environments [40]. In numerical simulations of a swarm of interacting monkey walks in a static but heterogeneous environment containing many impurities, the decay of memory was seen to improve the aggregation of individuals around the best resource sites, a phenomenon identified as a successful collective learning [41]. Very few analytical results are available on the relation between forgetting and spatial learning. To start with, is perfect memory a necessary condition for localization in preferential random walk models?
@@
We now present the structure of the article and summarize the main results. In Section II we introduce the memory random walk model with one impurity on the infinite lattice, where forgetting is incorporated via a general kernel. In Section III, we recall known analytical results on the steady state solution of the problem with perfect memory. Section IV reports simulation results. When memory has an asymptotic decay slower than $1/\tau$ (where $\tau$ is the time backward into the past), i.e., of the form $1/\tau^{\beta}$ with $0\leq\beta\leq 1$, we find that the system tends toward the same localized state at late times than with perfect memory or $\beta=0$ (Section IV.1). This fact is consistent with the prediction obtained from a decoupling approximation. However, with kernels decaying faster than $1/\tau$ (Sections IV.2 and IV.3), an unusual dynamical behavior is observed: localization becomes intermittent in time, or spatial learning is disrupted, i.e., well-localized periods of finite durations are separated by possibly very long intervals of diffusive motion. Somehow surprisingly, during the localized periods, the occupation probability of the lattice sites is very close to the solution with perfect memory. In Section V, we perform a linear stability analysis of the fully localized solution using the decoupling approximation, which turns out to be consistent only for memory kernels decaying more slowly than $1/\tau$. In these cases, we find that the localized solutions are stable, in agreement with the numerics of Section IV.1. We then study the relaxation at late times toward these stable states (Section VI), showing that it is non-exponential and of the form $t^{\nu}/\ln t$, with $\nu\in(-1,0)$ a non-trivial exponent solution of Eq. (56). In the boundary case $\beta=1$ corresponding to the $1/\tau$ memory kernel at the edge of intermittency, the localized state is still stable and the relaxation towards the latter is actually the fastest among all the values of $\beta\in[0,1]$: the amplitudes of the perturbations tend to zero no more slowly than $(\ln t)^{\alpha}/t$, with $\alpha$ an exponent [Eqs. (61)-(62)]. This speed-up, confirmed by numerical simulations, is opposite to the usual critical slowing down for a state that is marginally stable. Hence, the $1/\tau$ forgetting law is in some sense optimal, as $(i)$ it achieves the same localization as perfect memory by being the most economical on memory, and $(ii)$ it localizes the fastest. In Section VII we discuss different possible scenarios for intermittent localization (or imperfect learning) and conclude in Section VIII.
@@ II The model
We revisit the discrete time random walk model with preferential relocation to previously visited sites, in the presence of one impurity [20]. The original model considers a $1d$ lattice and a random walker with perfect memory that performs symmetric n.n. random steps with probability 1q1-q, and resets to a previously visited site with probability $q$. In the latter case, a site is chosen, among the visited sites, with a probability proportional to the total amount of time spent on that site since $t=0$. Additionally, an impurity site is located at the origin $n=0$, representing a food resource site. When the walker, with position $X_{t}$ at time $t$, is located at the origin (or $X_{t}=0$), it stays on that site for the next time-step with probability $\gamma$ and moves according to the two movement rules above with the complementary probability $1-\gamma$. Hence, the parameter $0\leq\gamma\leq 1$ represents the attractiveness of the origin, which is occupied for a longer time at each visit compared to the other sites of the lattice. This site is likely to be more strongly reinforced by the resetting dynamics.
As mentioned above, when memory is used during the time-step $t\to t+1$, a previous site is chosen to be revisited with a probability proportional to the accumulated amount of time spent there. This linear rule is equivalent to choosing a time $t^{\prime}$ in the past, $t^{\prime}=0,1,...,t$, with uniform probability and to relocate to the position occupied at time $t^{\prime}$, i.e., to set
@@
A site that has been often occupied will have more possibilities (times $t^{\prime}$) of being chosen. The fact that $t^{\prime}$ is distributed uniformly also means that the walker remembers equally well all its previous positions, or memory does not decay.
@@
Let us denote as $P_{n}(t)$ the walker’s occupation probability of the site $n$ at time $t$, $P_{n}(t)={\rm Prob}[X_{t}=n]$, given an unspecified initial condition. This quantity satisfies the following master equation,
@@
The first line of Eq. (II) represents the Markov random walk steps and trapping by the impurity, where $\delta_{i,j}$ is the Kronecker function. The second line accounts for the probability of revisiting the site $n$ via memory at time $t+1$ (because it was visited at a previous time $t^{\prime}$) by jumping from the origin which is occupied at time $t$. The third line counts the relocation jumps to $n$ from another site than the origin. The key point in this problem is that the resetting probability actually depends on the position: it is $q(1-\gamma)$ at the origin (since the walker does not move with probability $\gamma$) and $q$ at the other sites. Such heterogeneity is crucial for the existence of localized solutions, but it also makes the problem very difficult to solve, as Eq. (II) cannot be reduced to a closed form involving the marginal distribution $P_{n}(t)$ only. Clearly, the master equation explicitly depends on higher-order joint probabilities such as ${\rm Prob}[X_{t^{\prime}}=n\ {\rm and}\ X_{t}=0]$, therefore a whole hierarchy should be written for the multiple-time distribution functions [21].
@@ III Previous results for the perfect memory case
We summarize the results obtained from the uniform memory kernel $\pi_{t,t^{\prime}}=1/(t+1)$, corresponding to no memory decay.
III.1 Homogeneous lattice ($\gamma=0$)
For completeness, let us first recall the main properties of the monkey walk on a homogeneous lattice free of impurity, corresponding to $\gamma=0$ [10, 35, 17]. This problem is more tractable since Eq. (II) closes in that case and can be recast as,
$P_{n}(t+1)=\frac{1-q}{2}\left[P_{n+1}(t)+P_{n-1}(t)\right]+\frac{q}{t+1}\sum_{t^{\prime}=0}^{t}P(n,t^{\prime})\,.$ (5)
This equation is linear and the mean-square displacement of a particle starting from the origin, $M_{2}(t)\equiv\sum_{n=-\infty}^{\infty}n^{2}P(n,t)$, can be obtained exactly [10],
$M_{2}(t)=\frac{1-q}{q}\sum_{k=1}^{t}\frac{1-(1-q)^{k}}{k}\simeq\frac{1-q}{q}\left[\ln(qt)\,+\,\gamma+{\rm h.o.t.}\right],$ (6)
where $\gamma$ is the Euler constant and h.o.t. represents higher order terms, decaying as $1/t$ at large $t$. Hence, diffusion is logarithmically slow in time but unbounded. At large $t$, the distribution $P_{n}(t)$ obeys a scaling form,
$P_{n}(t)\simeq\frac{1}{\phi(t)}G\left(\frac{n}{\phi(t)}\right)\,$, (7)
with $\phi(t)=\sqrt{M_{2}(t)}$ and $G(z)=\frac{1}{\sqrt{2\pi}}e^{-\frac{z^{2}}{2}}$ the centered Gaussian distribution of unit variance. The occupation probability in Eq. (7) is therefore always time-dependent, tends to 0 uniformly at late times and no localized solutions exist.
III.2 Localization with $\gamma\neq 0$
In this case, let us assume that $P_{n}(t)$ in Eq. (II) tends to a stationary distribution $P_{n}^{(st)}>0$ at large times for all $n$, peaked around the origin [20]. This solution, if it exists, corresponds to a localized phase. Some progress can be made in the resolution of Eq. (II) with $\gamma\neq 0$ by noticing that the times $t^{\prime}$ and $t$ are in general very far apart asymptotically, justifying the use of a de-correlation approximation,